OpenStudy (loser66):

show that if A is a linear transformation such that A^2 +2A +3 = 0, then A is invertible Please, help

OpenStudy (tkhunny):

What are the properties of a linear transformation?

OpenStudy (loser66):

you mean definition?

OpenStudy (loser66):

It is just have product 's property like left/right distribute apply on addition and multiplication, nothing special

OpenStudy (tkhunny):

And there is a zero property.

OpenStudy (loser66):

yes

OpenStudy (tkhunny):

And for invertibility one must have in inverse transformation f(g(v)) = v AND g(f(u)) = u.

OpenStudy (loser66):

I am sorry for being dummy. I know that thing but don't see the link between that and the problem

OpenStudy (anonymous):

are all square matrices linear transformation matrices?

OpenStudy (loser66):

yes, invertible or not is another topic

OpenStudy (tkhunny):

?? Why? Are we just ignoring the problem statement? "...then A is invertible".

OpenStudy (anonymous):

this equation can be factored into (A-r1 I)(A-r2 I)=0, yielding 2 trivial invertible solution. i'm not quite sure how to prove these are the only possible solutions though

OpenStudy (loser66):

is it not that we have to simplify the left hand side to have the form of A* something = 1 to get that something is A inverse. then conclude that A is invertible

OpenStudy (tkhunny):

Do we agree that A MUST be square?

OpenStudy (loser66):

I have a simpler problem like A^2 -A +1 =0 --> A -A^2 =1 --> A(1-A) =1 and conclude that 1-A is inverse of A

OpenStudy (anonymous):

that makes sense. A^2+2A+4I=I factor & solve

OpenStudy (loser66):

and one more back ward to get A is inverse of (1-A) ... so that A is invertible

OpenStudy (loser66):

@eashy yyyyyyyyyyyyyyyyyyes!! I am so tired with it to not know how to square it

OpenStudy (anonymous):

solve the quadratic equation x^2+2x+4=0

OpenStudy (loser66):

oh my god.

OpenStudy (loser66):

Thanks a lot.

OpenStudy (anonymous):

:)

OpenStudy (loser66):

ha!! no no , it's not that , just A^2 +2A +1 =-2 or (A+1)^2 = -2

OpenStudy (anonymous):

\(A^2+2A=-3\Rightarrow -\frac{1}{3}A^2-\frac{2}{3}A=I\Rightarrow A\left(-\frac{1}{3}A-\frac{2}{3}I\right)=I\Rightarrow -\frac{1}{3}A-\frac{2}{3}I=A^{-1}\)

OpenStudy (anonymous):

\(-\frac{1}{3}A-\frac{2}{3}I=A^{-1}\) BTW, \(A\in \{n\times n \text{ matrices with complex elements, }n \ge 1\}\) otherwise \(A^2\) is undefined.